There were 3482 Walmart stores int he US in 2007 and 3808 Walmart store in the US in 2011. Wirte a liner equation in slope intercepts form that models this growth. Let x stand for the number of years after 2007 and y stand for the number of walmart stores in the US
show all work as you solve the linear modeling problem

Answer by jorel1380(3719)

3808-3482=326
2011-2007=4
326/4=81.5
So, the equation of the line would be:
y=3482+81.5x, where x equals the number of years after 2007.
☺☺☺☺